[Brahms-l] Shift report 20040309 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 09 2004 - 07:42:30 EST
------------------------------------------------------------
Date: Tue Mar  9 07:42:29 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 80 - 80

Shift Summary:
------------------------------------------------------------
1.  WE had three stores.  One of them had poor ZDC rates so we, the RHIC experimenters, voted to dump such that we can get luminosity until CAD begins with their work.

2. I took a lot of data and a tpc and a non tpc pedestal run.

3. I didn't experience any detector problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 9 07:42:36 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 09 2004 - 07:42:56 EST